OpenStudy (anonymous):

if the height of a refrigerator is 6 feet and the base is 3 feet. find the angle of inclination. a. 67.43 b. 26.56 c. 63.43

OpenStudy (anonymous):

 to find the angle B, you have to find the hypotenuse, which is \[\sqrt{45}\] then you can use any of the trigonometric functions. i prefer sin to find the angle you start with: \[\sin \theta= opp/hyp\] look at the drawing. the opposite side is 6, therefore: \[\sin \theta=6/\sqrt{45}\] use the inverse function of sin \[\theta=\sin^{-1} (6/\sqrt{45})\] and you get \[\theta \approx 63.43\]